OpenStudy (anonymous):

a) Find all fourth roots of 1 in polar form. b) Express them in Cartesian form. c) Show how they can be expressed as powers of one fixed fourth root of 1.

OpenStudy (amistre64):

4th roots in polar form divide the unit circle into 4 equal parts.... so 90 degree seperations

OpenStudy (amistre64):

since the any root of 1 = 1; im assuming they want (1,1) (1,-1) (-1,-1) (-1,1) as answers

OpenStudy (anonymous):

isnt it x^4=1?

OpenStudy (amistre64):

I was thinking \(\sqrt[4]{x}\) at \(x=1\)

OpenStudy (anonymous):

but i guess it is x=1^(1/4) this is what we call four forth root of 1

OpenStudy (anonymous):

x=1,-1,i,-i

OpenStudy (amistre64):

90s or the 45s....

OpenStudy (amistre64):

if we go with the i stuff; its the 90s and thats prolly the better interpretatio

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

but for 1 its 0

OpenStudy (anonymous):

for -1, pi

OpenStudy (amistre64):

there is no pi in the cartesian; just your 1s and 0s for your intercepts

OpenStudy (amistre64):

(1,0) (0,1) (-1,0) (0,-1)

OpenStudy (amistre64):

but what 'c' is asking for I dunno

OpenStudy (amistre64):

maybe \((1-0i)^{1/4}\) ?

OpenStudy (amistre64):

or is it simply \(i^4\)

OpenStudy (anonymous):

i cant guess any idea

OpenStudy (anonymous):

i^4 seems better

OpenStudy (amistre64):

im thinking the first since that implies a complex plane and 4 roots

OpenStudy (anonymous):

but how (-1,0) and(0,-1)

OpenStudy (amistre64):

\((1+0i)^4\) maybe? if forget if its ^4 or ^(1/4) that pops out 4 times

OpenStudy (amistre64):

\(sqrt{-6}\) has complex roots right?

OpenStudy (anonymous):

look...when we talk about the cube roots of 1 , how we express it? x=1^1/3..no?

OpenStudy (amistre64):

i believe so

OpenStudy (anonymous):

n for forth root it is x=1^1/4

OpenStudy (anonymous):

\[x^4 = 1\]

OpenStudy (amistre64):

sqrt(-9) = 3i and we can find both those roots in the complex plane ... gonna have to dbl chk with the wolfram :)

OpenStudy (anonymous):

\[(x^2-1)(x^2+1)=0\]

OpenStudy (anonymous):

x=1,-1,i,-i

OpenStudy (anonymous):

for x=1 r=1 and theta =0

OpenStudy (anonymous):

x=-1 r=1, theta =pi -1=cospi

OpenStudy (anonymous):

4th roots of 1 are 1, -1, i, -1

OpenStudy (anonymous):

you know one answer is 1. divide unit circle (in complex plane) into 4 equal parts and you will see i, -1, -i

OpenStudy (anonymous):

did u get part c?

OpenStudy (anonymous):

what is part c?

OpenStudy (anonymous):

read the post:P

OpenStudy (anonymous):

oh yes they are all powers of i

OpenStudy (anonymous):

i, i^2, i^3, i^4 finito

OpenStudy (anonymous):

no question says" Show how they can be expressed as powers of one fixed fourth root of 1."

OpenStudy (anonymous):

that fixed root is i.

OpenStudy (anonymous):

you cannot express i as a power of 1. i assume they mean integral powers